OpenStudy (calculusxy):

MEDAL! Two cards are selected from a deck of ten card from 1-10. Once a card is selected, it is not replaced. What is the probability that the first card selected is less than 5 and the second card is 5 or more?

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

\[\frac{ 5 }{ 10 } \times \frac{ 5 }{ 9 }\]

OpenStudy (calculusxy):

Is that the correct expression?

jimthompson5910 (jim_thompson5910):

"What is the probability that the first card selected is less than 5" that's 4/10 since there are 4 numbers (1,2,3,4) that are less than 5 out of 10 total

OpenStudy (calculusxy):

\[\frac{ 4 }{ 10 } \times \frac{ 5 }{ 9 }\]

jimthompson5910 (jim_thompson5910):

"and the second card is 5 or more" the numbers 5,6,7,8,9,10 are 5 or more. There are 6 of these numbers out of 10-1 = 9 total

OpenStudy (calculusxy):

So my expression is correct right?

OpenStudy (calculusxy):

It would be 20/90 = 2/9

jimthompson5910 (jim_thompson5910):

it should be (4/10)*(6/9)

OpenStudy (calculusxy):

Why 6?

OpenStudy (calculusxy):

Yes

jimthompson5910 (jim_thompson5910):

the numbers 5,6,7,8,9,10 are 5 or more. There are 6 of these numbers

OpenStudy (calculusxy):

Uh...I felt like I forgot to count :(

OpenStudy (calculusxy):

so: 4/10 x 6/9

OpenStudy (calculusxy):

That's correct right?

jimthompson5910 (jim_thompson5910):

correct, you'd multiply and reduce